OpenStudy (anonymous):

Jack mows his lawn in 3 hrs. Marilyn can mow the same yard in 4 hrs. How much time would it take them to mow the yard together?

OpenStudy (anonymous):

((1/3)+(1/4))x=1 where x is the total time he can do a third of the lawn per hour and she can only do a quarter per hour so it works out to be around 1.741hours by the above equation

OpenStudy (anonymous):

t/3 + t/4 = 1...solve for t......t=12/7 hrs